# HiFICL: High-Fidelity In-Context Learning for Multimodal Tasks

### 13 Mar 2026 v1

@@ 5 Conclusion
Furthermore, our work contributes a new perspective to Parameter-Efficient Fine-Tuning (PEFT). By instantiating the hypothesis of ICL as ”inference-time finetuning,” HiFICL introduces a novel form of dynamic, content-aware adaptation. We believe this principle of context-aware parameterization charts a promising new frontier for developing more expressive and interpretable PEFT technologies.
@@ 7.3 Additional Qualitative Examples
We provide qualitative examples on VQAv2 and COCO in Fig. 8 to complement our quantitative results. The examples, generated using the LLaVA-Interleave-7b model, visually demonstrate HiFICL’s ability to produce more faithful responses. For instance, HiFICL often corrects factual errors made by other methods (e.g., identifying a ”fire engine” instead of a ”car”) and reduces object hallucination (e.g., avoiding the erroneous ”parking meters”), showcasing its effectiveness in capturing nuanced visual details.
@@ 8.1 Overview of Recent LoRA Variants
LoRA-style PEFT has evolved rapidly beyond its original formulation. As illustrated in Figure 9, a significant body of work has focused on innovating the structure of the low-rank update matrix, $\Delta W$. Representative variants include methods that employ Mixture-of-Experts (MoE) principles to dynamically select different update matrices (MoE-based LoRA) or introduce structured sparsity for higher efficiency (FlyLoRA). Despite their architectural diversity, these methods share a common foundational principle: they all operate in the weight space by applying an additive update to the model’s pre-trained matrices.
HiFICL can be viewed as a paradigm shift in this landscape. Instead of innovating how the $\Delta W$ matrix is constructed, HiFICL innovates on where and how the adaptation is applied. It moves the adaptation from the static weight space to the dynamic activation space, directly modeling the ICL-induced shift. Interestingly, because HiFICL also employs learnable, low-rank matrices (our virtual key-value pairs), the structural innovations from the LoRA family are conceptually orthogonal and potentially complementary to our approach. For instance, one could envision a future variant, ”MoE-HiFICL,” that learns multiple sets of virtual key-value pairs and dynamically routes to them. This highlights that HiFICL not only provides a standalone solution but also opens a new, parallel branch for the future development of dynamic, context-aware PEFT.

### 26 Mar 2026 v3

@@ 5 Conclusion
Beyond ICL approximation, our work bridges the gap with Parameter-Efficient Fine-Tuning (PEFT). By practically instantiating the concept of “inference-time finetuning,” HiFICL demonstrates the viability of dynamic, content-aware adaptation. We hope this context-aware parameterization approach can serve as a useful reference for future developments in efficient model tuning.
Acknowledgment
This work was supported by the National Natural Science Foundation of China under Grant No. 62472072.
@@ 7.3 Additional Qualitative Examples
We provide qualitative examples on VQAv2 and COCO in Fig. 8 to complement our quantitative results. The examples, generated using the LLaVA-Interleave-7b model, visually demonstrate HiFICL’s ability to produce more faithful responses. For instance, HiFICL often corrects factual errors made by other methods (e.g., identifying a “fire engine” instead of a “car”) and reduces object hallucination (e.g., avoiding the erroneous “parking meters”), showcasing its effectiveness in capturing nuanced visual details.
@@ 8.1 Overview of Recent LoRA Variants
LoRA-style PEFT has evolved rapidly. As shown in Figure 9, variants like MoE-based LoRA or FlyLoRA innovate the low-rank update matrix ($\Delta W$). However, they all operate in the static weight space via additive updates.
HiFICL shifts this paradigm: instead of modifying the weight space, it operates in the dynamic activation space to directly model the ICL-induced shift. Because HiFICL also employs low-rank virtual matrices, existing LoRA structural innovations are potentially complementary. For instance, a future “MoE-HiFICL” could dynamically route to multiple virtual key-value pairs, opening new avenues for context-aware PEFT.
@@
8.3 Performance at Optimal Rank ($r=8$)
As discussed in the main text, the default rank setting ($r=16$) inherited from prior ICL approximation baselines can lead to overfitting for LoRA on certain datasets. To provide a generic and fair PEFT comparison, we evaluate both LoRA and HiFICL using an optimized rank of $r=8$, which empirically serves as a robust “sweet spot” for both methods.
[…]
Table 8: Performance comparison between LoRA and HiFICL with an optimal rank of $r=8$. Best results for each model are bolded.
Model Method # Params (M) VQAv2 OK-VQA COCO
LLaVA LoRA 9.8 ($\times$4.5) 75.75 54.28 1.3307
HiFICL 2.2 ($\times$1.0) 74.66 53.19 1.3315
Idefics2 LoRA 8.8 ($\times$4.0) 69.58 60.18 1.3448
HiFICL 2.2 ($\times$1.0) 72.08 58.96 1.2851